Move It, Move Na! riders receive a fixed pay of Php30 per delivery. On 

top of that, they receive a 2% commission based on the amount of food 

orders. Rider A was able to have five food deliveries with the following 

bill amounts: Php 275.00; 560.00; 138.00; 629.50; and 1,750.00. How 

much will Rider A receive for his daily compensation? If you are the 

manager, will you also be basing the riders’ compensation on a per 

delivery basis? Why?


1
Expert's answer
2021-12-13T10:09:03-0500

Rider A

Fixed pay 30

Commissions  

0.02*275 = 5.5

0.02*560 = 11.2

0.02*138 = 2.76

0.02*629.5 = 12.59

0.02*1750 = 35

Total 97.05 phps


As the manager I will also base the commission on per delivery basis as doing so will motivate the rider to do more deliveries per day much faster